Determine the pH of the following solution:

0.954 M NaCl
Answer #1

Answer : pH of 0.954 M NaCl solution = 7

Since NaCl is a salt formed by strong acid ( HCl) and strong base (NaOH) .

NaCl + H​​​​​​2​​​​​O ----> Na+ (aq) + Cl​​​​​- (aq)

No further Hydrolysis of any ion formed .

So, pH will be only due to dissociation of water .

H​​​​​​2​​​​​O ----> H+ (aq) + OH​​​​​- (aq)

From the dissociation of water , Kw = 10-14 = [H+ ] [OH​​​​​- ]

So, [H+ ] = 10-7 = [OH​​​​​- ]

pH = -log(H+ ) = -log(10-7) = 7